OpenStudy (anonymous):

Algebra 2 question. Will using f(g(x)) and g(f(x)) result in the same number? If I use them both for y=2x+15.

OpenStudy (anonymous):

I need to finish this in the next 5-10 minss

OpenStudy (anonymous):

" Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will result in the same number. "

OpenStudy (anonymous):

f(g(x)) is not always equal g(f(x)) f(x) = x^2 + 2x g(x) = x^3 - 1 so f(g(2)) = f(8-1) = f(7) = 49 + 14 = 63 and g(f(2)) = g(8) = 4096 - 1 = 4095 and is not equal ;)

OpenStudy (anonymous):

can you help me do it for y=2x+15 I don't really get it.

OpenStudy (anonymous):

@mhmdrz91

OpenStudy (anonymous):

what is f(x) and what is g(x)? you just say y=2x+15!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

they didnt give me

OpenStudy (anonymous):

did u read the directions up there that i wrote? ^^

OpenStudy (anonymous):

" Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will result in the same number. "

OpenStudy (anonymous):

i guess i have to make up something?

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

maybe ;) where this question discussed?

OpenStudy (anonymous):

its my homework

OpenStudy (debbieg):

Does it give you a function somewhere for f(x) and g(x)? Or maybe a function for f(x), and ask you to find g(x)=inverse of f(x)?

OpenStudy (jdoe0001):

hmmm @cupcake111 I think they're referring to the inverse of f(x), that is the inverse of 2x+15 do they result in the same thing, provided that g(x) is the inverse of f(x) poor wording but I think it's what's meant

OpenStudy (jdoe0001):

in which case, yes \(\bf f(f^{-1}(x)) = x, \qquad f^{-1}(f(x)) = x\)

OpenStudy (jdoe0001):

assuming \(\bf g(x) = f^{-1}(x)\)

OpenStudy (anonymous):

no it has nothing to do with inverse

OpenStudy (anonymous):

this is what i got so far

OpenStudy (jdoe0001):

so, what does it mean? :S

OpenStudy (anonymous):

f(x)= 2x+15 if x=5

OpenStudy (anonymous):

f(x)= 2(5)+15 f(x)= 25

OpenStudy (anonymous):

f(x) --> y= 2x+15 f(g) ->> 25

OpenStudy (anonymous):

and then i dont know

OpenStudy (jdoe0001):

hmmm, where's the function g(x)? what is it equal to? I can see f(x) = 2x+5, but what about g(x)?

OpenStudy (anonymous):

i haven't thought of anything

OpenStudy (anonymous):

btw, its 2x+15, not 2x+5

OpenStudy (debbieg):

You can't evaluate anything involving g(x) if g(x) isn't defined anywhere. Are you sure it isn't somewhere near the top of the problem set, something like that?

OpenStudy (debbieg):

You posted above that the problem says: " Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will result in the same number. " Where did you find the equation you are using for f(x)? It isn't in that statement.

OpenStudy (jdoe0001):

based on the wording, g(x) is most likely the inverse of f(x)

OpenStudy (jdoe0001):

it's just poorly worded really

OpenStudy (anonymous):

its from one of the previous questions @DebbieG

OpenStudy (debbieg):

That's what I think too. I wonder if somewhere in this problem set, you had to find g(x)=inverse of f(x)??

OpenStudy (anonymous):

there are 6 questions in total, this is the last one, it is a lot of words but i will put it so u can see and understand better

OpenStudy (debbieg):

OK, then what do the previous questions say about g(x)?? LOL

OpenStudy (anonymous):

no i didnt

OpenStudy (jdoe0001):

heehhe

OpenStudy (jdoe0001):

lack of clairvoyant abilities put us in the precarious position to have to ask you, what's g(x)?

OpenStudy (anonymous):

the charts got messed up

OpenStudy (jdoe0001):

hmmm, take a screenshot :|

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

Wow. Badly, badly worded. I would ask your teacher what g(x) is, or are you supposed to make one up.

OpenStudy (anonymous):

she is not replying by text or email :(

OpenStudy (anonymous):

i think you make one up

OpenStudy (anonymous):

my mom doesn't even get it! sigh

OpenStudy (anonymous):

i still have more hw... and only an hour and a half

OpenStudy (debbieg):

Well, honestly, I think she wants you to make up f(x) (because really, it doesn't even SAY that f(x) is your made-up function, but I think it's implied) and then I suspect she wants g(x) to be the inverse of f(x). But again, it's really not spelled out in the problem.

OpenStudy (anonymous):

:/

OpenStudy (debbieg):

I would just use g(x)=f{inverse}(x). And then say something like, "Assuming that g(x)=f{inverse}(x)...." and proceed from there.

OpenStudy (jdoe0001):

\( \text{5. Using complete sentences, describe to the Martians}\\ \\ \text{how to find the} \color{red}{\text{ inverse of your function}}\)

OpenStudy (anonymous):

i've already done that

OpenStudy (anonymous):

5. To find the inverse of the function, the first step is to replace the f(x) with y. F(x)= 2x+15 will be y= 2x+15. Then, switch the x and y variables. You will have x=2y+15. Next, solve for y by isolating it on one side of the equal sign. x-15= 2y+15-15 x-15= 2y x-15 divided by 2 = y x-15 over 2 = f-1(x)

OpenStudy (jdoe0001):

then step 6 is to show that \(\bf f(f^{-1}(x)) = x, \qquad f^{-1}((x)) = x\\ g(x) = f^{-1}(x)\)

OpenStudy (debbieg):

Looks good.

OpenStudy (anonymous):

how exactly..?

OpenStudy (debbieg):

just evaluate f(g(x)) and g(f(x)). Plug your g(x) into f for x, and then plug your g(x) into g for x.

OpenStudy (debbieg):

Simplify... and you should get x each time.

OpenStudy (anonymous):

but i dont know my g(x)

OpenStudy (debbieg):

Sure you do... you just explained to the martians how to cook it up! LOL....

OpenStudy (jdoe0001):

cupcake111 i've already done that <----

OpenStudy (anonymous):

i know..

OpenStudy (jdoe0001):

well, so the inverse of 2x+15,IS g(x)

OpenStudy (debbieg):

To find the inverse of the function, the first step is to replace the f(x) with y. F(x)= 2x+15 will be y= 2x+15. Then, switch the x and y variables. You will have x=2y+15. Next, solve for y by isolating it on one side of the equal sign. x-15= 2y+15-15 x-15= 2y x-15 divided by 2 = y x-15 over 2 = f-1(x) \[g(x)=f ^{-1}(x)=\frac{ x-15 }{ 2 }\]and \[f(x)=2x+15\]

OpenStudy (jdoe0001):

\(\bf f(x)= \color{red}{2x+15}\\ g(x) = \color{blue}{\cfrac{x-15}{2}}\\ f(g(x)) = \color{red}{2\left(\color{blue}{\cfrac{x-15}{2}}\right)+15}\\ g(f(x)) = \color{blue}{\cfrac{(\color{red}{2x+15})-15}{2}}\)

OpenStudy (jdoe0001):

so simplify each, see if they give the same value

OpenStudy (anonymous):

Ok thank u so much!! let me just do it..

OpenStudy (anonymous):

both of u :))

OpenStudy (debbieg):

You're welcome.

OpenStudy (anonymous):

wait but the parenthesis for the f(g(x)) one, those big parenthesis, how do i get rid of those

OpenStudy (debbieg):

just multiply. :) The 2 on the outside is just a product of 2 x the big fraction inside, so what will happen when you do that?

OpenStudy (debbieg):

Something will conveniently cancel. :)

OpenStudy (anonymous):

i cant do it!!! helpp and u can't cancel it outtt

OpenStudy (anonymous):

@DebbieG